3) Assume that two independent investments opportunities, X and Y, exist in the market. Investment provides an expected return of 12% with a standard deviation of 12%, while investment Y yields 10% return with a standard deviation of 10%.

a. Calculate the expected return for a portfolio that consists of 40% of X and 60% of Y.

b. What is the standard deviation of such portfolio?
Answer #1

a)
Expected return = 0.4*12 + 0.6*10 = 10.8 %

b)
std. dev. = sqrt(0.4^2*12^2 + 0.6^2*10^2)
= 7.68 %